# Pay for The Second-Best Service: A Game-Theoretic Approach Against Dishonest LLM Providers

### v3

@@ A.1. Impossibility of First-Best User Utility
Definition 3 (Type and Type Set).
@@
Assume there exists an $o(T)$-approximate incentive-compatible mechanism that guarantees the user a worst-case utility of $\xi u_{FB}+(1-\xi)u_{SB}$. Denote $i^{*}$ as the provider who delivers $u_{FB}$, and $i^{\prime}$ as the provider who delivers $u_{SB}$. We construct $N>\exp\left(\frac{1}{\xi-\lambda}\right)$ type sets, where $\lambda=\frac{\xi}{2}$, and for the $j$-th type set, we set $u_{FB}=T\cdot(\frac{1}{2}+\frac{j}{2N})$ and $u_{SB}=\frac{T}{2}$. we denote $u^{(j)}=\frac{1}{2}+\frac{j}{2N}$ and $u^{\circ}=\frac{1}{2}$. Within these type sets, we consistently use the notation $(i^{*},i^{\prime})$. The types of all other service providers are identical across all type sets. Let $F_{j}(c)=\frac{cu^{(j)}}{p_{i^{*}}}$, where =1=1 and $p_{i^{*}}$ keeps unchanged for all type sets.
@@
$\Leftrightarrow x_{j}-\frac{y_{j}}{u^{(j)}}$ $>x_{j^{\prime}}-\frac{y_{j^{\prime}}}{u^{(j)}}-\frac{\lambda}{u^{(j)}N}$
@@
$=\left(\xi-\frac{j\lambda}{N}\right)\left(1+\frac{1}{j}\sum_{l=1}^{j-1}\frac{j-l}{l}\right)$
@@
$=\left(\xi-\frac{j\lambda}{N}\right)\left(1+\sum_{l=1}^{j-1}\frac{1}{l}-\frac{j-1}{j}\right)$
@@ A.2. Properties of Our Mechanism
We first state all lemmas used in the following proofs.
Lemma A.1 (Hoeffding’s inequality).
Let $Z_{1},\ldots,Z_{n}$ be independent bounded random variables with $Z_{i}\in[a,b]$ for all $i$, where $\infty<a\leq b<\infty$. Then
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\left(Z_{i}-\mathbb{E}\left[Z_{i}\right]\right)\geq t\right)\leq\exp\left(-\frac{2nt^{2}}{(b-a)^{2}}\right)$
and
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\left(Z_{i}-\mathbb{E}\left[Z_{i}\right]\right)\leq-t\right)\leq\exp\left(-\frac{2nt^{2}}{(b-a)^{2}}\right)$
for all $t\geq 0$.
Lemma A.2 (Union bound).
For a countable set of events $A_{1},A_{2},A_{3},\ldots$, we have
$\mathbb{P}\left(\bigcup_{i=1}^{\infty}A_{i}\right)\leq\sum_{i=1}^{\infty}\mathbb{P}\left(A_{i}\right).$
@@
Proof.
The total completed queries
[…]
The second-to-last inequality is from the definition of $\delta_{i}$ in Mechanism 1.
@@
On the other hand, the objective function $H(c,l)=p_{i}l-cl$ has bounded partial derivatives on the compact feasible domain, which implies it is also Lipschitz continuous. This property guarantees that a perturbation of order $O(M)$ in the input variables will result in a change of at most order $O(M)$ in the function’s value,i.e.
@@
See 5.2 Next we give the proof of Theorem 5.2, we do not repeat it here due to lack of space.
Proof.
@@
This implies that the provider will not trigger the validated flag (turning it to false) with a probability of $1-o\left(\frac{1}{T}\right)$. Therefore, the expected utility for provider $i^{*}$ in the exploitation phase is then given by
$U_{2}\geq T_{R}(p_{i^{*}}l^{\prime}-c^{\prime}l^{\prime}),$
where “2” in the notation $U_{2}$ refers to the second phase.
@@
Then it incurs a cost of $c^{\prime\prime}$ and generates a corresponding token sequence $\tau^{\prime\prime}$. If $|\tau^{\prime\prime}|\leq l^{\prime\prime}$, it reports an output token sequence of length $l^{\prime\prime}$; otherwise, it reports the truthful token sequence. Should the provider be sufficiently fortunate, this strategy would also allow it to avoid triggering the validated flag and thereby achieve an ideal utility. Although the realization of such a scenario is probabilistically negligible, the resulting ideal utility serves as a definitive upper bound for the true optimal utility. Consequently, this ideal utility can be leveraged to compute a more strict additive approximation ratio. Hence, we calculate the ideal provider utility
$U_{2}^{\prime}=T_{R}(p_{i^{*}}l^{\prime\prime}-c^{\prime\prime}l^{\prime\prime}).$
and the additive approximation ratio
$U_{2}^{\prime}-U_{2}$ $=T_{R}[p_{i^{*}}(l^{\prime\prime}-l^{\prime})-(c^{\prime\prime}l^{\prime\prime}-c^{\prime}l^{\prime})].$

### v4

@@ A.1. Impossibility of First-Best User Utility
For clarity, we first provide definitions of types and type sets.
Definition 3 (Types and Type Sets).
@@
Assume there exists an $o(T)$-approximate incentive-compatible mechanism that guarantees the user a worst-case utility of $\xi u_{FB}+(1-\xi)u_{SB}$. Denote $i^{*}$ as the provider who delivers $u_{FB}$, and $i^{\prime}$ as the provider who delivers $u_{SB}$. We construct $N>\exp\left(\frac{1}{\xi-\lambda}\right)$ type sets, where $\lambda=\frac{\xi}{2}$, and for the $j$-th type set, we set $u_{FB}=T\cdot(\frac{1}{2}+\frac{j}{2N})$ and $u_{SB}=\frac{T}{2}$. we denote $u^{(j)}=\frac{1}{2}+\frac{j}{2N}$ and $u^{\circ}=\frac{1}{2}$. Within these type sets, we consistently use the notation $(i^{*},i^{\prime})$. The types of all other service providers are identical across all type sets. Let $F_{j}(c)=\frac{cu^{(j)}}{p_{i^{*}}}$, where $p_{i^{*}}$ keeps unchanged for all type sets.
@@ A.2. Properties of Our Mechanism
We first state two lemmas used in the following proofs.
@@
We omit this proof due to lack of space.
@@
On the other hand, the objective function $H(c,l)=p_{i}l-cl$ has bounded partial derivatives on the compact feasible domain, which implies it is also Lipschitz continuous. This property guarantees that a perturbation of order $O(M)$ in the input variables will result in a change of at most order $O(M)$ in the function’s value, i.e.,
@@
Next we give the proof of Theorem 5.2, we do not repeat it here due to lack of space.
Proof of Theorem 5.2.
@@
This implies that the provider will not trigger the validated flag (turning it to false) with a probability of $1-o\left(\frac{1}{T}\right)$. Therefore, the expected utility for provider $i^{*}$ in the exploitation phase is then given by $U_{2}\geq T_{R}(p_{i^{*}}l^{\prime}-c^{\prime}l^{\prime})$, where “2” in the notation $U_{2}$ refers to the second phase.
@@
Then it incurs a cost of $c^{\prime\prime}$ and generates a corresponding token sequence $\tau^{\prime\prime}$. If $|\tau^{\prime\prime}|\leq l^{\prime\prime}$, it reports an output token sequence of length $l^{\prime\prime}$; otherwise, it reports the truthful token sequence. Should the provider be sufficiently fortunate, this strategy would also allow it to avoid triggering the validated flag and thereby achieve an ideal utility. Although the realization of such a scenario is probabilistically negligible, the resulting ideal utility serves as a definitive upper bound for the true optimal utility. Consequently, this ideal utility can be leveraged to compute a more strict additive approximation ratio. Hence, we know the ideal provider utility $U_{2}^{\prime}=T_{R}(p_{i^{*}}l^{\prime\prime}-c^{\prime\prime}l^{\prime\prime})$, and the additive approximation ratio $U_{2}^{\prime}-U_{2}=T_{R}[p_{i^{*}}(l^{\prime\prime}-l^{\prime})-(c^{\prime\prime}l^{\prime\prime}-c^{\prime}l^{\prime})]$.